OpenStudy (anonymous):

Solve 4^(2x) = 7^(x−1).

OpenStudy (anonymous):

@Astrophysics @rishavraj @ganeshie8

OpenStudy (freckles):

have you tried taking ln( ) of both sides

OpenStudy (anonymous):

wouldn't it be log?

OpenStudy (freckles):

\[\ln(4^{2x})=\ln(7^{x-1}) \\ \\ \text{ Now use power rule } \\ 2x \ln(4)=(x-1) \ln(7)\] you can use log too if you want

OpenStudy (freckles):

not distribute on the right hand side

OpenStudy (freckles):

and put your terms with x in it on one side and your terms without x on the opposing side

OpenStudy (anonymous):

1.38629436112(2x) I mean, sorry :)

OpenStudy (freckles):

are you allowed to approximate ?

OpenStudy (anonymous):

no, not until we have the final answer

OpenStudy (freckles):

ok so leave that one side as 2ln(4)x and distribute on the other side recall the distributive property is a(b+c)=ab+ac

OpenStudy (freckles):

use distributive property here and what do you get: \[\ln(7)(x-1)=?\]

OpenStudy (anonymous):

1.94591014906

OpenStudy (anonymous):

1.94591014906x-1.94591014906

OpenStudy (freckles):

I think you mean to say ln(7)x-ln(7)

OpenStudy (anonymous):

oh yes! Sorry

OpenStudy (freckles):

\[\ln(4^{2x})=\ln(7^{x-1}) \\ \\ \text{ Now use power rule } \\ 2x \ln(4)=(x-1) \ln(7) \\ 2 \ln(4) x=\ln(7) x- \ln(7) \]

OpenStudy (freckles):

now to get the x people together subtract ln(7)x on both sides

OpenStudy (freckles):

\[ax=bx-c \\ ax-bx=-c \\ x(a-b)=-c \\ \text{ the last step is to choose to divide by } \\ \text{ what is in front of the } x \\ \\ \text{ in this example that would be } (a-b) \\ x=\frac{-c}{a-b}\]

OpenStudy (freckles):

this is exactly what you are going to do here first subtract ln(7)x on both sides then use my example as a guide sorta to finally find x

OpenStudy (anonymous):

2.77258872224x=1.94591014906x-1.94591014906

OpenStudy (freckles):

I thought you aren't allowed to approximate until you have final answer

OpenStudy (anonymous):

i didn't :)

OpenStudy (freckles):

ln(7) is irrational there is no way you can write out the whole number same for ln(4)

OpenStudy (anonymous):

oh. Sorry, I just typed it into my calculator & thats what it said

OpenStudy (freckles):

your calculator can only show you so many digits

OpenStudy (freckles):

it doesn't have an infinite screen display

OpenStudy (freckles):

anyways... \[\ln(4^{2x})=\ln(7^{x-1}) \\ \\ \text{ Now use power rule } \\ 2x \ln(4)=(x-1) \ln(7) \\ 2 \ln(4) x=\ln(7) x- \ln(7) \\ \text{ the step I was asking you \to do is subtract } \ln(7) x \text{ on both sides } \\ 2 \ln(4) x -\ln(7) x=-\ln(7)\] try factoring the x out on the left hand side like I did in my example ax-bx=x(a-b)

OpenStudy (anonymous):

2 In(x)-In(-3)x=-In (7)

OpenStudy (freckles):